Victor’s is open from 11:30am until 9:00pm 5 days per week, which works out to 250 days per year considering holiday closures. Victor recently invested in Toast, a mobile POS system for taking orders and issuing checks. The new system, which cost $35,000, provides key information on sales and costs, while also facilitating the job of employees who take orders and issue the checks. Using this new system, Victor obtained the following information on his results for 2019:

Sales from Food: $430,000 (excludes tips)
Sales from Drinks: $240,000 (excludes tips)
Number of customer bills/checks 18,000 (includes drinks and food)
Direct Costs of Food: $200,000
Direct Costs of Drinks: $ 80,000
Wages for hourly Workers: $140,000 (includes all employment costs)
Salary of Manager: $ 60,000 (includes all employment costs)
Other operating Expenses $ 80,000 (rent, insurance, supplies, real estate taxes, etc.)

Please calculate the following for Victor’s Bar and Grill:
a) Average bill for food: ______
b) Average bill for drinks: ________
c) Average CM for food: _________ (CM = contribution margin)
d) Average CM for drinks: ______
e) CM % on food: _______
f) CM % on drinks: _______
g) Total CM %: _______
h) Total Revenue: _______
i) Total CM: ________
j) Operating Expenses: _______
k) Operating income: ________
l) Average hourly labor cost per bill: _______

In 2020, Victor’s expects the following changes to his business:
a) Wages for workers will increase by 15% because of a new minimum wage.
b) He will increase the salary of the manager by 5%.
c) All other costs, including material costs per meal, are not expected to change.
d) He plans to increase his prices of food by 3% and to increase his drink prices by 5%.
e) The number of customers (meals) is expected to increase by 2%

Using this new information, please forecast the following for 2020:
a) Average bill for food: ______
b) Average bill for drinks: ________
c) Average CM for food: _______
d) Average CM for drinks: ______
e) Average total CM per bill: ________
f) Total CM %: _______
g) Total Revenue: _______
h) Total CM: ________
i) Operating income: ________
j) Average hourly labor cost per bill: _______
Answer #1

For 2019

a) $430,000/18,000 = $23.88/bill (average bill for food)

b) $240,000/18,000 = $13.33/bill (average bill for drinks)

c) Contribution from food = $430,000-$200,000 = $230,000

Thus, Average CM from food = $230,000/18,000 = $12.77/bill

d) Contribution from drinks = $240,000-$80,000 = $160,000

Thus, Average CM from drinks = $160,000/18,000 = $8.88/bill

e) CM% on food = $12.77*100/$23.88 = 53.47%

f) CM% on drinks = $8.88*100/$13.33 = 66.61%

g) Total CM = $230,000+$160,000 = $390,000

Total Sale = $430,000+$240,000 = $670,000

Thus, Total CM% = $390,000*100/$670,000 = 58.20%

h) Total Revenue = $430,000+$240,000 = $670,000

i) Total CM = $230,000+$160,000 = $390,000

j) Operating expenses = $ (200,000+80,000+140,000+60,000+80,000) = $ 560,000

k) Opearting Income = $ (430,000+240,000) = $670,000

l) Avg labour cost = $140,000/(250 days * 9.5 hrs/day) = $58.94/hour

For 2020

New wages = $140,000*1.15 = $161,000

New salary = $60,000*1.05 = $63,000

New No. of customers = 18,000*1.02 = 18,360

New Sale : Food = ($430,000*18,360/18,000)*1.03 = $451,758

Drinks = ($240,000*18,360/18,000)*1.05 = $257,040

a) Avg food bill = $451,758/18,360 = $24.60/bill

b) Avg drinks bill = $257,040/18,360 = $14/bill

c) Contribution from food = $451,758-$200,000 = $251,758

Thus, Average CM from food = $251,758/18,360 = $13.71/bill

d) Contribution from drinks = $257,040-$80,000 = $177,040

Thus, Average CM from drinks = $177,040/18,360 = $9.64/bill

e) Average Total Contribution = $ (451,758+257,040)/18,360 = $38.60/bill

f) Total CM = $251,758+$177,040 = $428,798

Total Sale = $451,758+$257,040 = $708,798

Thus, Total CM% = $428,798*100/$708,798 = 60.50%

g) Total Revenue = $451,758+S257,040 = $708,798

h) Total CM = $251,758+$177,040 = $428,798

i) Opearting Income = $451,758+$257,040 = $708,798

j) Avg labour cost = $161,000/(250 days * 9.5 hrs/day) = $67.79/hour